OpenStudy (anonymous):

how do you work 1,240 in to prime factorization and list two other multile that equal 1240

OpenStudy (anonymous):

Keep dividing it by prime numbers until you end up with a prime number. 1240 = 31 X 5 X 2 X 2 X 2. 1240 is also a multiple of 31, 310, 20, 62, etc...

OpenStudy (anonymous):

Well it looks _really_ even. So I'd start by looking at powers of 2. \[1240/2^2 = 310\] still even...\[1240/2^3 = 155\]Not even. So ok, that's all the 2's\[155/5 = 31\]that's all the 5's and 31 is prime. So we have:\[1240 = 2^3\times 5 \times 31\]

OpenStudy (anonymous):

thanks